7 L of an acid solution was mixed with 3 L of a 15% acid solution to make a 29% acid solution. Find the percent concentration of the first solution.

9 lbs. of mixed nuts containing 55% peanuts were mixed with 6 lbs. of another kind of mixed nuts that contain 40% peanuts. What percent of the new mixture is peanuts?
A solution containing 30% chicken broth is to be mixed with a solution containing 50% chicken broth to make 200L containing 42% chicken broth. How much of each solution should be mixed?

House blend coffee is 50% Columbian beans and Special blend coffee is 80% Columbian beans. How much of each should be used to produce 100kg of a blend that is 68% Columbian beans?

How many quarts of a 50% sugar solution must be mixed with 8 quarts of water to make a 10% solution?

Answer by stanbon(75887)

7 L of an acid solution was mixed with 3 L of a 15% acid solution to make a 29% acid solution. Find the percent concentration of the first solution.
acid + acid = acid
(x/100)*7 + 0.15*3 = 0.29*(7+3)
-------
Multiply thru by 100 to get:
7x + 45 = 290
7x = 245
x = 35
---
Ans:: 35%
-----------------------------
9 lbs. of mixed nuts containing 55% peanuts were mixed with 6 lbs. of another kind of mixed nuts that contain 40% peanuts. What percent of the new mixture is peanuts?
peanuts + peanuts = peanuts
0.55*9 + 0.40*6 = x(9+6)
Multiply thru by 100 to get
55*9 + 40*6 = 1500x
x = 49%
--------------------
Comment:: The three problems below can be worked in a similar manner.
Cheers,
Stan H.
------------

A solution containing 30% chicken broth is to be mixed with a solution containing 50% chicken broth to make 200L containing 42% chicken broth. How much of each solution should be mixed?
House blend coffee is 50% Columbian beans and Special blend coffee is 80% Columbian beans. How much of each should be used to produce 100kg of a blend that is 68% Columbian beans?
How many quarts of a 50% sugar solution must be mixed with 8 quarts of water to make a 10% solution?